OpenStudy (anonymous):

Find two numbers whose sum is 29 and whose difference is 15

OpenStudy (diyadiya):

x+y=29 x-y=15

OpenStudy (mr.math):

Let the numbers be x and y, then: \(x+y=29\) \(x-y=15\) Add the two equations you get 2x=44.. You can take it from here.

OpenStudy (diyadiya):

Do elimination method or substitution method

OpenStudy (anonymous):

22 and 7

OpenStudy (anonymous):

Alright, we see! ^^

OpenStudy (anonymous):

Thanks you guys!

OpenStudy (anonymous):

So, we have 2x=44 /2 /2 ________ x= 22 44-22=7 x=22 and y=7

OpenStudy (diyadiya):

Yes

OpenStudy (anonymous):

Sweet! Thanks again! ^^